# LEGO-Eval: Towards Fine-grained Evaluation on Synthesizing 3D Embodied Environments with Tool Augmentation

### v1

@@ 1 Introduction
Embodied agents represent a paradigm shift from digital assistants to physical collaborators (shridhar2020alfred; brohan2022rt; chang2025partnr). While training agents in the physical world is feasible, it is impractical due to the slow pace of real-time learning and the escalating costs of scaling across multiple environments. Consequently, training in realistic simulators has become the dominant approach, allowing agents to learn real-world physics and semantics through navigation and interaction in 3D scenes (xiang2020sapien; makoviychuk2isaac; li2022igibson). Training in realistic scenes is critical, as unrealistic scenes can prevent agents from learning the physical and semantic understanding of scene components—such as objects, walls, doors, windows, and rooms—ultimately limiting their effectiveness in real-world deployment. For instance, an agent trained in a kitchen without a refrigerator may fail to locate one, open it, and retrieve items during deployment, due to insufficient understanding of the presence and function of the refrigerator. While simulators aim to provide such realism, popular platforms such as AI2THOR (kolve2017ai2) rely on manually created 3D environments by experts—a labor-intensive and costly process that severely limits diversity. This limitation in number of scenes can lead to failures in real-world tasks that involve much greater variation in environment.
@@
To scale up the number of scenes, recent work has explored automatic scene generation using Large Language Models (LLMs) and random sampling (deitke2022; yang2024holodeck). These methods often generate scenes without explicit textual guidance, or rely only on coarse-grained instructions (e.g., “Modern-style kitchen”). While effective for generating many scenes, they frequently yield unrealistic environments—such as a kitchen missing a refrigerator or bookshelves obstructing a window. Training in these scenes can mislead agents to develop incorrect understandings of the physical world. To mitigate this, scene generation can be guided by fine-grained instructions that describe the real-world environment in detail. Thus, fully satisfying the constraints in such detailed instructions is crucial for generating realistic scenes that enable robust agent training.
However, existing evaluation methods cannot reliably verify constraint satisfactions. This unreliability stems from their inability to perform multi-hop grounding: (1) identifying scene components mentioned in the instruction, and (2) verifying their attributes and spatial relationships. For example, assessing the constraint “a blue chair placed next to the black desk” involves locating both objects in the scene, verifying their colors, and evaluating their spatial relationship. Current approaches lack capability for such multi-hop grounding. One widely adopted method, CLIPScore (hessel2021clipscore), is inadequate due to the limited capacity of CLIP (radford2021learning) to interpret complex 3D scenes (hegde2023clip; ma2023crepe). Similarly, vision-language models (VLMs) used as judges also struggle with precise localization of scene components, as shown in Figure 1, which limits their ability to assess attributes or spatial relations (li2024topviewrs).
@@ 2 Related Work
Early work explored rule-based pipelines, spatial priors, and database composition (coyne2001wordseye; chang2014learning; ma2018language). With the advent of neural generative models, diffusion-based pipelines have emerged as a dominant paradigm for text-to-3D synthesis (hollein2023text2room; ma2024fastscene; tang2024diffuscene; zhou2025roomcraft; fang2025ctrl). Recently, LLMs and VLMs have also been adopted for indoor scene generation, leveraging their real-world priors to act as effective compositional planners. Prior works generally fall into three categories. Scene-level generation approaches (feng2023layoutgpt; yang2024holodeck; bucher2025respace) generate complete 3D environments directly from language instructions. Object selection and placement methods (yang2024llplace; ccelen2025design) focus on selecting relevant objects and determining their placements. Meanwhile, layout optimization methods (sun2025layoutvlm; ran2025direct) aim to optimize the spatial arrangement of predefined assets.
@@
Evaluating text-guided 3D scene synthesis with fine-grained instructions requires verifying whether generated scene includes all specified constraints, which is a complex task involving multi-hop grounding. For example, assessing “a blue chair placed next to the black desk” demands identifying objects, verifying attributes, and checking spatial relationships. Current methods typically rely on two approaches: (1) CLIPScore (ocal2024sceneteller; fu2024anyhome; deng2025global), which measures the CLIP similarity between a top-down image of the scene and the instruction, (2) prompting VLMs (wang2024architect; ccelen2025design; sun2025layoutvlm; ling2025scenethesis), providing VLMs with the instruction along with images of the scene captured from multiple viewpoints. However, these methods lack deep understanding of 3D scenes, which limits multi-hop grounding (ma2023crepe; li2024topviewrs). To address this, SceneEval (tam2025sceneeval) predefines evaluation criteria such as object count and attributes, and object-object, object-architecture spatial relations, but cannot evaluate attributes and spatial relations of architectures (e.g., “Sliding window is on orange wall”). The spatial relations are also predefined to basic relations (e.g., left, near), limiting evaluation of diverse spatial expressions such as “The table is closer to the chair than bed”. In contrast, LEGO-Eval supports spatial reasoning across all scene components and can handle a broad range of relationships expressed in natural language.
@@
LLMs are often augmented with external tools to overcome its limitations of parametric memory. Many works augment LLMs with external tools to improve the aspects such as factuality (komeili2022internet), math abilities (imani2023mathprompter), and solve complex tasks (paranjape2023art; chen2023chatcot; patil2024gorilla). There has also been growing interest in augmenting language models with multimodal tool use. VisProg (gupta2023visual) and ViperGPT (suris2023vipergpt) integrate vision modules with Python-based text utilities to decompose and solve visual reasoning tasks. In contrast, AVIS (hu2023avis) and Chameleon (lu2023chameleon) combine vision tools with textual functions such as knowledge retrieval or tabular processing, targeting broader multimodal information-seeking and compositional reasoning. Our evaluation method similarly augments VLMs with multimodal tools for environment interaction, textual reasoning, and multimodal reasoning, supporting multi-hop grounding in 3D scenes.
@@ 3.1 Evaluation Framework
Step 1: Constraint Identification. To enable rigorous evaluation, we begin by identifying constraints $C=(c_{1},\dots,c_{k})$ within each instruction $I$. Fine-grained instructions include multiple constraints, each contributing to aspects of the desired scene. LEGO-Eval identifies these constraints and categorizes them into one of four types, similar to the modules in Holodeck (yang2024holodeck):
@@ 3.2 Tool Set
  • Multimodal Reasoning: These tools convert visual information into textual descriptions to retrieve specific information from images. Since VLMs struggle with multi-image inputs (wang2024muirbench), we use LLMs and VLMs to convert images into text descriptions.
@@ 4.1.1 Setup
Baselines. We compare our performance with SceneEval (tam2025sceneeval), CLIPScore (hessel2021clipscore), and VLM-as-a-judge. CLIPScore performs binary judgment on the instruction and top-down scene image using thresholds of 15, 20, and 25. For VLM-as-a-judge, we provide scene images from four perspectives and use Gemini-2.5-Pro, GPT-4o mini, and GPT-4.1 as base models with self-consistency across 3 samples for fair comparison. Following LEGO-Eval, all baseline methods consider a scene aligned with the instruction if all constraints in the instruction are satisfied.
@@ 4.1.2 Results
In Table 1, we present a comparative evaluation of LEGO-Eval against existing methods. SceneEval (tam2025sceneeval), constrained by a fixed set of metrics and criteria, is unable to assess 41% of the constraints in the dataset. To enable a fair comparison, we evaluate its performance under two settings: treating unevaluable constraints as incorrect, and modifying instructions to include only those constraints it can process. SceneEval also suffers from a fundamental limitation—it fails to reliably ground scene components referenced in the instructions. While accurate evaluation depends first on identifying the relevant objects, SceneEval often fails at this initial step.
Predominantly used methods exhibit similar limitations. CLIPScore (hessel2021clipscore) is unreliable for fine-grained alignment with 3D scenes, as CLIP (radford2021learning) is trained on 2D image–text pairs and lacks 3D spatial understanding. VLM-as-a-judge also frequently misidentifies or fails to localize mentioned components, resulting in incorrect constraint assessments. In contrast, LEGO-Eval achieves higher F1 and Cohen’s kappa scores at both the constraint and instruction levels. By integrating diverse tools for multi-hop grounding, LEGO-Eval more effectively locates and interprets scene components—enabling accurate and robust evaluation of 3D scene synthesis.
@@ 4.2.1 Setup
Baselines. We evaluate four LLM-based 3D scene synthesis methods on LEGO-Bench: LayoutGPT (feng2023layoutgpt), Holodeck (yang2024holodeck), I-Design (ccelen2025design), and LayoutVLM (sun2025layoutvlm). These methods vary in functionality: Holodeck generates complete scenes with object selection, attributes, and placement; I-Design selects and places objects; LayoutGPT and LayoutVLM position a given set of objects. To enable fair comparison, we augment the latter three with Holodeck to produce full scenes for evaluation.
@@ 4.2.2 Results
We benchmark LLM-based 3D scene synthesis methods on LEGO-Bench to provide a comprehensive assessment of existing approaches. As shown in Table 3, most methods achieve average partial SRs exceeding 50%, yet their performance on object selection and placement is lower. This indicates that current methods struggle to effectively handle these fundamental aspects of scene synthesis. Interestingly, while selecting appropriate objects affects the performance of object placement, we also observe vice versa. Although LayoutGPT (feng2023layoutgpt) and LayoutVLM use Holodeck (yang2024holodeck) for object selection, their performance differs from that of Holodeck. This stems from differences in how these methods position the selected objects, with some methods failing to place them in the scene.
@@ 5 Analysis
LEGO-Eval as a feedback signal for refinement. To illustrate the reliability and interpretability of evaluations, we refine the results of Holodeck (yang2024holodeck) from Section 4.2 with the feedback from LEGO-Eval. Specifically, we use LEGO-Eval to evaluate scenes generated by Holodeck, then provide feedback to refine invalid scenes. We iterate this refinement process 3 times, comparing LEGO-Eval against prompted VLMs as feedback signals. The results show that Holodeck achieves higher performance using LEGO-Eval as feedback than prompted VLMs, demonstrating LEGO-Eval’s superior feedback quality for refinement. A possible explanation is that while prompted VLMs often follow invalid reasoning paths, LEGO-Eval uses valid reasoning, which is reflected in its detailed, interpretable evaluation explanations.
@@ B.2 Details of Human Annotation
In the third phase, annotators create a scene that fully validates the constraints within the instruction. To achieve this, we first input the instructions into Holodeck (yang2024holodeck), which generates scenes in textual representations (i.e., JSON scripts) that can be rendered in the 3D simulator. The textual representations contain comprehensive scene information, including the asset ID of each component as well as the exact position and rotation of all scene components. As shown in Table 4.2, Holodeck cannot generate scenes that fully align with the instructions. Thus, annotators manually modify the text representation of the scene so that the scene fully aligns with the instruction.
@@ C.1 Model Inference
We prompt the closed-source models, including Gemini 2.5 Pro, o4-mini, gpt-4.1-2025-04-14, and gpt-4.1-mini-2025-04-14, for both LLM and VLM inference. The input images are resized to 1200px resolution, and the temperature is fixed at 0 to ensure deterministic outputs.
For the open-source model Gemma3-27B, Qwen2.5VL-32B, and Qwen3-32B, we use the default hyperparameter settings provided by the model for inference. The input images are resized to 335px resolution; when the number of images exceeds the model’s processing capacity, two images are concatenated into a single input, with their corresponding names displayed at the top-left corner of each image.
@@ Appendix F Limitations
Appendix G Examples of Tool Image Outputs
@@
Appendix H Prompts used in our Works

### v2

@@ 1 Introduction
Embodied agents represent a paradigm shift from digital assistants to physical collaborators (Shridhar et al., 2020; Brohan et al., 2022; Chang et al., 2025). While training agents in the physical world is feasible, it is impractical due to the slow pace of real-time learning and the escalating costs of scaling across multiple environments. Consequently, training in realistic simulators has become the dominant approach, allowing agents to learn real-world physics and semantics through navigation and interaction in 3D scenes (Xiang et al., 2020; Makoviychuk et al., 2021; Li et al., 2022). Training in realistic scenes is critical, as unrealistic scenes can prevent agents from learning the physical and semantic understanding of scene components—such as objects, walls, doors, windows, and rooms—ultimately limiting their effectiveness in real-world deployment. For instance, an agent trained in a kitchen without a refrigerator may fail to locate one, open it, and retrieve items during deployment, due to insufficient understanding of the presence and function of the refrigerator. While simulators aim to provide such realism, popular platforms such as AI2THOR (Kolve et al., 2017) rely on manually created 3D environments by experts—a labor-intensive and costly process that severely limits diversity. This limitation in number of scenes can lead to failures in real-world tasks that involve much greater variation in environment.
@@
To scale up the number of scenes, recent work has explored automatic scene generation using Large Language Models (LLMs) and random sampling (Deitke et al., 2022; Yang et al., 2024b). These methods often generate scenes without explicit textual guidance, or rely only on coarse-grained instructions (e.g., “Modern-style kitchen”). While effective for generating many scenes, they frequently yield unrealistic environments—such as a kitchen missing a refrigerator or bookshelves obstructing a window. Training in these scenes can mislead agents to develop incorrect understandings of the physical world. To mitigate this, scene generation can be guided by fine-grained instructions that describe the real-world environment in detail. Thus, fully satisfying the constraints in such detailed instructions is crucial for generating realistic scenes that enable robust agent training.
However, existing evaluation methods cannot reliably verify constraint satisfactions. This unreliability stems from their inability to perform multi-hop grounding: (1) identifying scene components mentioned in the instruction, and (2) verifying their attributes and spatial relationships. For example, assessing the constraint “a blue chair placed next to the black desk” involves locating both objects in the scene, verifying their colors, and evaluating their spatial relationship. Current approaches lack capability for such multi-hop grounding. One widely adopted method, CLIPScore (Hessel et al., 2021), is inadequate due to the limited capacity of CLIP (Radford et al., 2021) to interpret complex 3D scenes (Hegde et al., 2023; Ma et al., 2023). Similarly, vision-language models (VLMs) used as judges also struggle with precise localization of scene components, as shown in Figure 1, which limits their ability to assess attributes or spatial relations (Li et al., 2024).
@@ 2 Related Work
Early work explored rule-based pipelines, spatial priors, and database composition (Coyne and Sproat, 2001; Chang et al., 2014; Ma et al., 2018). With the advent of neural generative models, diffusion-based pipelines have emerged as a dominant paradigm for text-to-3D synthesis (Höllein et al., 2023; Ma et al., 2024; Tang et al., 2024; Zhou et al., 2025; Fang et al., 2025). Recently, LLMs and VLMs have also been adopted for indoor scene generation, leveraging their real-world priors to act as effective compositional planners. Prior works generally fall into three categories. Scene-level generation approaches (Feng et al., 2023; Yang et al., 2024b; Bucher and Armeni, 2025) generate complete 3D environments directly from language instructions. Object selection and placement methods (Yang et al., 2024a; Çelen et al., 2025) focus on selecting relevant objects and determining their placements. Meanwhile, layout optimization methods (Sun et al., 2025; Ran et al., 2025) aim to optimize the spatial arrangement of predefined assets.
@@
Evaluating text-guided 3D scene synthesis with fine-grained instructions requires verifying whether generated scene includes all specified constraints, which is a complex task involving multi-hop grounding. For example, assessing “a blue chair placed next to the black desk” demands identifying objects, verifying attributes, and checking spatial relationships. Current methods typically rely on two approaches: (1) CLIPScore (Öcal et al., 2024; Fu et al., 2024; Deng et al., 2025), which measures the CLIP similarity between a top-down image of the scene and the instruction, (2) prompting VLMs (Wang et al., 2024b; Çelen et al., 2025; Sun et al., 2025; Ling et al., 2025), providing VLMs with the instruction along with images of the scene captured from multiple viewpoints. However, these methods lack deep understanding of 3D scenes, which limits multi-hop grounding (Ma et al., 2023; Li et al., 2024). To address this, SceneEval (Tam et al., 2025) predefines evaluation criteria such as object count and attributes, and object-object, object-architecture spatial relations, but cannot evaluate attributes and spatial relations of architectures (e.g., “Sliding window is on orange wall”). The spatial relations are also predefined to basic relations (e.g., left, near), limiting evaluation of diverse spatial expressions such as “The table is closer to the chair than bed”. In contrast, LEGO-Eval supports spatial reasoning across all scene components and can handle a broad range of relationships expressed in natural language.
@@
LLMs are often augmented with external tools to overcome its limitations of parametric memory. Many works augment LLMs with external tools to improve the aspects such as factuality (Komeili et al., 2022), math abilities (Imani et al., 2023), and solve complex tasks (Paranjape et al., 2023; Chen et al., 2023; Patil et al., 2024). There has also been growing interest in augmenting language models with multimodal tool use. VisProg (Gupta and Kembhavi, 2023) and ViperGPT (Surís et al., 2023) integrate vision modules with Python-based text utilities to decompose and solve visual reasoning tasks. In contrast, AVIS (Hu et al., 2023) and Chameleon (Lu et al., 2023) combine vision tools with textual functions such as knowledge retrieval or tabular processing, targeting broader multimodal information-seeking and compositional reasoning. Our evaluation method similarly augments VLMs with multimodal tools for environment interaction, textual reasoning, and multimodal reasoning, supporting multi-hop grounding in 3D scenes.
@@ 3.1 Evaluation Framework
Step 1: Constraint Identification. To enable rigorous evaluation, we begin by identifying constraints $C=(c_{1},\dots,c_{k})$ within each instruction $I$. Fine-grained instructions include multiple constraints, each contributing to aspects of the desired scene. LEGO-Eval identifies these constraints and categorizes them into one of four types, similar to the modules in Holodeck (Yang et al., 2024b):
@@ 3.2 Tool Set
  • Multimodal Reasoning: These tools convert visual information into textual descriptions to retrieve specific information from images. Since VLMs struggle with multi-image inputs (Wang et al., 2024a), we use LLMs and VLMs to convert images into text descriptions.
@@ 4.1.1 Setup
Baselines. We compare our performance with SceneEval (Tam et al., 2025), CLIPScore (Hessel et al., 2021), and VLM-as-a-judge. CLIPScore performs binary judgment on the instruction and top-down scene image using thresholds of 15, 20, and 25. For VLM-as-a-judge, we provide scene images from four perspectives and use Gemini-2.5-Pro, GPT-4o mini, and GPT-4.1 as base models with self-consistency across 3 samples for fair comparison. Following LEGO-Eval, all baseline methods consider a scene aligned with the instruction if all constraints in the instruction are satisfied.
@@ 4.1.2 Results
In Table 1, we present a comparative evaluation of LEGO-Eval against existing methods. SceneEval (Tam et al., 2025), constrained by a fixed set of metrics and criteria, is unable to assess 41% of the constraints in the dataset. To enable a fair comparison, we evaluate its performance under two settings: treating unevaluable constraints as incorrect, and modifying instructions to include only those constraints it can process. SceneEval also suffers from a fundamental limitation—it fails to reliably ground scene components referenced in the instructions. While accurate evaluation depends first on identifying the relevant objects, SceneEval often fails at this initial step.
Predominantly used methods exhibit similar limitations. CLIPScore (Hessel et al., 2021) is unreliable for fine-grained alignment with 3D scenes, as CLIP (Radford et al., 2021) is trained on 2D image–text pairs and lacks 3D spatial understanding. VLM-as-a-judge also frequently misidentifies or fails to localize mentioned components, resulting in incorrect constraint assessments. In contrast, LEGO-Eval achieves higher F1 and Cohen’s kappa scores at both the constraint and instruction levels. By integrating diverse tools for multi-hop grounding, LEGO-Eval more effectively locates and interprets scene components—enabling accurate and robust evaluation of 3D scene synthesis.
@@ 4.2.1 Setup
Baselines. We evaluate four LLM-based 3D scene synthesis methods on LEGO-Bench: LayoutGPT (Feng et al., 2023), Holodeck (Yang et al., 2024b), I-Design (Çelen et al., 2025), and LayoutVLM (Sun et al., 2025). These methods vary in functionality: Holodeck generates complete scenes with object selection, attributes, and placement; I-Design selects and places objects; LayoutGPT and LayoutVLM position a given set of objects. To enable fair comparison, we augment the latter three with Holodeck to produce full scenes for evaluation.
@@ 4.2.2 Results
We benchmark LLM-based 3D scene synthesis methods on LEGO-Bench to provide a comprehensive assessment of existing approaches. As shown in Table 3, most methods achieve average partial SRs exceeding 50%, yet their performance on object selection and placement is lower. This indicates that current methods struggle to effectively handle these fundamental aspects of scene synthesis. Interestingly, while selecting appropriate objects affects the performance of object placement, we also observe vice versa. Although LayoutGPT (Feng et al., 2023) and LayoutVLM use Holodeck (Yang et al., 2024b) for object selection, their performance differs from that of Holodeck. This stems from differences in how these methods position the selected objects, with some methods failing to place them in the scene.
@@ 5 Analysis
LEGO-Eval as a feedback signal for refinement. To illustrate the reliability and interpretability of evaluations, we refine the results of Holodeck (Yang et al., 2024b) from Section 4.2 with the feedback from LEGO-Eval. Specifically, we use LEGO-Eval to evaluate scenes generated by Holodeck, then provide feedback to refine invalid scenes. We iterate this refinement process 3 times, comparing LEGO-Eval against VLM-as-a-Judge as feedback signals. The results show that Holodeck achieves higher performance using LEGO-Eval as feedback than VLM-as-a-Judge, demonstrating LEGO-Eval’s superior feedback quality for refinement. A possible explanation is that while prompted VLMs often follow invalid reasoning paths, LEGO-Eval uses valid reasoning, which is reflected in its detailed, interpretable evaluation explanations.
@@ B.2 Details of Human Annotation
In the third phase, annotators create a scene that fully validates the constraints within the instruction. To achieve this, we first input the instructions into Holodeck (Yang et al., 2024b), which generates scenes in textual representations (i.e., JSON scripts) that can be rendered in the 3D simulator. The textual representations contain comprehensive scene information, including the asset ID of each component as well as the exact position and rotation of all scene components. As shown in Table 4.2, Holodeck cannot generate scenes that fully align with the instructions. Thus, annotators manually modify the text representation of the scene so that the scene fully aligns with the instruction.
@@ C.1 Model Inference
We prompt the closed-source models, including Gemini 2.5 Pro, o4-mini, gpt-4.1-2025-04-14, and gpt-4.1-mini-2025-04-14, for both LLM and VLM inference. The input images are resized to 1200px resolution, and we use the default hyperparameter settings provided by the model for inference.
For the open-source model Gemma3-27B, Qwen2.5VL-32B, and Qwen3-32B, the temperature is fixed at 0 to ensure deterministic outputs. The input images are resized to 335px resolution; when the number of images exceeds the model’s processing capacity, two images are concatenated into a single input, with their corresponding names displayed at the top-left corner of each image.
@@ Appendix F Limitations
Appendix G Acknowledgments
This work utilized the AI2-THOR simulator (Kolve et al., 2017) as the primary environment for conducting experiments. We further acknowledge Objaverse (Deitke et al., 2023) for providing large-scale, high-quality 3D assets that enabled the construction of diverse and realistic indoor scenes. These publicly available resources have significantly supported research in 3D scene generation leveraging LLMs.
Appendix H Examples of Tool Image Outputs
@@
Appendix I Prompts used in our Works